# Evolutionary Prediction Games

### v2

@@ 5 Avenues for Coexistence
Figure 3: Graphical illustration of theoretical coexistence results. Top row presents data distributions, bottom row shows the corresponding prediction games. (Left) Stable and mutualistic coexistence induced by the use of a proxy loss (Theorem 1). (Center) Stable coexistence induced by interpolation (Section D.7). (Right) Mutualistic coexistence induced by a finite training set (Section D.8).
@@ 5.1 Learning in Practice
While the goal in learning is to minimize the expected 0-1 loss, in practice learning often resorts to optimizing proxy objectives over sampled data via empirical risk minimization. Here we show that, perhaps surprisingly, basic aspects of this approach—namely the use of surrogate losses, access to finite data, and interpolation methods—can create conditions that facilitate mutualism, in which all groups benefit from coexistence. For each aspect, we demonstrate through a carefully crafted example how it can act as a mechanism that enables stable coexistence. Here we focus our discussion on surrogate loss as a representative construction; for finite data and memorization, we illustrate the underlying distribution and induced game in Figure 3, and defer the formal analysis to Section D.6.
@@
The construction is outlined in Figure 3 (Left), and proof is provided in Section D.6. The proof leverages the bias of hinge loss against minority classes: For each group alone, minimizing the hinge loss biases the learned classifier against the minority class, which is suboptimal. But since each group has a different minority class, these biases cancel out at the mixed equilibrium. The surprising property is that this state is also evolutionarily stable: when one group grows slightly (i.e., $p_{1}=0.5+\epsilon$), it loses more from the fact that the other group shrinks (i.e., $p_{2}=0.5-\epsilon$) than it gains from its own growth. In this sense, each group needs the other to maintain its existence, giving rise to stable and mutualistic symbiosis. Section D.6.3 expands on this idea to show that varying the amount of regularization leads to bifurcations, in which the game transitions between 3, 1, and 5 equilibria.
@@
In the example game above, as in those we construct for finite data (Section D.8; Figure 3 (Right)) and interpolation (Section D.7; Figure 3 (Center)), mutualism is enabled by how user self-selection compensates for the algorithm’s imperfections. When the resulting bias is complementary across groups, this can give rise to (implicit) cooperation. One interpretation is that the existence of each group acts as regularization on others: from the perspective of group $k$, the objective can be written as $\operatorname*{arg\,max}_{h}\mathrm{acc}_{k}(h)+\lambda R(h)$ where $R(h)=\sum_{j\neq k}\mathrm{acc}_{k}^{\prime}(h)$ is a data-dependent regularizer with coefficient $\lambda=\frac{1}{p_{k}}\sum_{j\neq k}p_{j}$. Nonetheless, our results should not be taken to imply that proxy losses or finite data are a good means to achieve coexistence; rather, they highlight how mutualism can organically materialize when learning affects different groups differently. We explore this further empirically in Section 6.1.
@@ 5.2 Stabilizing Coexistence Equilibria
Here we focus on a particular aspect of this general question that arises from our last result in Sec. 4, namely: given a system with a desirable but unstable mixed equilibrium ${\bm{p}}^{*}$, how do we stabilize it? We propose a conceptually simple solution that works by reweighing examples in the standard accuracy objective in a way that accounts for dynamcis. The main idea is to invert the natural tendency of dynamics to push away from the unstable equilibrium, and instead pull towards it, by training ‘as if’ the distribution was at a different state ${\bm{p}}^{\prime}$ than its actual current state ${\bm{p}}$:
@@
Proof in Section D.9. The idea is that $\mathcal{A}^{\prime}$ trains ‘as if’ the state is $2{\bm{p}}^{*}-{\bm{p}}$, which can be achieved by weighing examples from each group $k$ by $w_{k}=\frac{p_{k}}{2p^{*}_{k}-p_{k}}$. This inverts the natural tendency of dynamics to push away from the unstable equilibrium, and instead pull towards it. The statement regards optimal classifiers and requires knowledge of ${\bm{p}}^{*}$, but in practice we can set $w$ according to an estimated $\tilde{{\bm{p}}}$. We further explore this approach in Section 6.2.
@@ 6 Experiments
Figure 4: Empirical evaluation results. (Left) Evolutionary prediction game induced by CIFAR-10 with Resnet-9 and groups representing horizontal flips. The game has unstable mutualism (Section 6.2). (Center) Simulated stabilization of replicator dynamics on the CIFAR-10 game (Section 5.2). (Right) Game induced by MNIST with label noise, showing stable mutualism (Section 6.1).
We now turn to explore evolutionary prediction games empirically using real data and simulated dynamics. Our goal here is to demonstrate how some of the principles we have discussed so far apply in more realistic settings. We include two experiments, designed to complement our results from sections 5.1 and 5.2, respectively. The first experiment shows how mutualism can arise organically due to group balancing, here from complementarities in label noise. The second demonstrates how stabilization can improve both overall and per-group accuracies by facilitating coexistence.
@@ 6.1 Stable Coexistence From Overparameterization
Modern neural networks are frequently overparameterized, and are known to achieve strong test-time performance despite memorizing the training data [e.g., 54]. Recent research attributes this phenomenon to the implicit biases introduced by the training process [e.g., 50]. Building on these ideas, here we demonstrate that stable coexistence can emerge when an overparameterized neural network is trained on data in which groups have complementing label noise. Our construction considers a multi-class settings where each group includes a majority of noisy examples from some class, and a minority of clean examples from other classes. This aims to captures settings in which user groups (e.g., by country) differ in the marginal distribution $p(y)$ over classes (e.g., spoken language), and where the dominant class is more likely to be misclassified (e.g., due to many accents).
@@
We use the MNIST dataset for digit recognition [30], which consists of 70,000 28x28 grayscale images of handwritten digits across 10 classes. Each class is split unevenly between the two groups: Group $A$ is biased towards even digits ${\left\{{0,2,4,6,8}\right\}}$, and group $B$ is biased towards odd digits, both with a 4:1 imbalance. For each group, we introduce label noise to the majority classes, mapping the true label of each digit $d$ to the next digit with same parity with probability $0.2$ (i.e. for group $A$, label noising stochastically maps $0\mapsto 2$, $2\mapsto 4$, etc.). We train a convolutional neural network for 200 epochs using stochastic gradient decent with momentum. The training and test sets are split independently, and for each ${\bm{p}}$ we measure the prediction model’s accuracy on both splits of the test sets (representing $\mathrm{acc}_{A}(h_{\bm{p}})$ and $\mathrm{acc}_{B}(h_{\bm{p}})$). Under the given parameters, the label noising process leads to an accuracy upper bound of $84\%$ assuming perfect recognition.
@@ Results.
Figure 4 (Right) shows the estimated evolutionary prediction game, averaged across 505050{}50 repetitions. The game has a stable coexistence equilibrium, with $80.4\pm 0.2\%{}$ test accuracy (vs. the theoretical upper bound of $84\%$ due to label noise). This indicates that groups naturally balance each other in this setting. Note that training accuracy is $98.7\pm 0.1\%{}$, suggesting that interpolation is indeed the likely phenomenon at play.
6.2 Mutualistic Coexistence From Data Augmentation
Data augmentation is a method for enriching the training set with class-preserving transformations of inputs (e.g., camera angle, source of lighting) [e.g., 46]. Augmentations are typically considered as artificial constructs, but there are also cases where they emerge naturally (e.g., medical images from different hospitals or imaging devices). Here we show that when different sources correspond to different groups, data pooling can serve to foster beneficial coexistence in which all groups gain from the existence of others. This suggests that augmentations remain effective in the long run.
@@ 6.2 Mutualistic Coexistence From Data Augmentation
We use the CIFAR-10 image recognition dataset [27], which consists of 60,000 32x32 color images in 10 classes, with 6,000 images per class. As horizontal image flips are considered class-preserving for this dataset, group $A$ consists of images sampled directly from CIFAR-10, and group $B$ consists of images under horizontal flip. We use a ResNet-9 network for prediction [20], and train it using the ffcv framework with default optimization parameters [29]. For each ${\bm{p}}$, we measure the prediction model’s accuracy on the original images from the CIFAR test set (representing $\mathrm{acc}_{A}(h_{\bm{p}})$), and on their flipped counterparts (representing $\mathrm{acc}_{B}(h_{\bm{p}})$). For stabilization, we use the method described in Section 5.2 with ${\bm{p}}^{*}=(0.5,0.5)$. We simulate the evolutionary process using discrete replicator dynamics, and use linear interpolation to determine intermediate fitness values.
@@ Results.
Figure 4 (Left) presented the estimated evolutionary prediction game, averaged across 202020{}20 repetitions. The game has three equilibria: Two stable equilibria with single-group dominance ($92.6\pm 0.1\%{}$ accuracy), and an unstable coexistence equilibrium ($93.5\pm 0.1\%{}$ accuracy). The mixed equilibrium is optimal for both groups in terms of accuracy, but requires stabilization to become feasible. Figure 4 (Center) shows how our algorithm from Section 5.2 is able to attain stable beneficial coexistence, despite using an empirical proxy objective.
@@ Acknowledgements
Appendix A Broader Impact and Limitations
@@
The bulk of our results are theoretical in nature, and as such, rely on several assumptions. Some regard the learning process: for example, we consider only simple accuracy-maximizing algorithms that optimize over a fixed-size training set sampled only from the current (and not past) data distribution. Others concern dynamics: for example, we assume no exogenous forces (i.e., which could prevent extinction), intra-group distribution shifts, or inter-group dependencies (other than those indirectly formed through the classifier). Whether our results hold also when these assumptions are relaxed is an important future question.
@@ B.1 Performative Prediction
Performative prediction [39] is a framework for modeling the effects of predictive models on the data distributions they are tested and trained on (see [18] for a recent survey). Dynamics induced by evolutionary prediction games are a special case of the stateful performative prediction framework [3, 41], which models stateful feedback loops. The performative effects of a predictor $h\in\mathcal{H}$ are formally modeled using a transition map $D_{t+1}=\mathcal{T}(D_{t},h)$, where $D_{t},D_{t+1}\in{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$ are the data distributions at times $t,t+1$, and $h\in\mathcal{H}$ is the predictor being deployed. The fixed points of the dynamics induced by evolutionary prediction games (i.e. equilibrium states ${\bm{p}}^{*}$), correspond to fixed point distributions as defined in [3, Definition 4], as they satisfy $D_{{\bm{p}}^{*}}=\mathcal{T}(D_{{\bm{p}}^{*}},h_{{\bm{p}}^{*}})$, where $h_{{\bm{p}}^{*}}$ is the classifier trained on $D_{{\bm{p}}^{*}}$. Moreover, we note that the notion of evolutionary stability is stronger than the notion of fixed point distributions, as evolutionarily stable points also remain stable under small perturbations around ${\bm{p}}^{*}$.
@@
Formally, given a collection of group distributions $D_{1},\dots,D_{K}\in{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$, denote the space of mixtures by $\Delta_{\mathrm{mix}}={\left\{{\sum_{k}p_{k}D_{k}\mid{\bm{p}}\in\Delta^{K}}\right\}}$, and note that $\Delta_{\mathrm{mix}}\subseteq{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$. Any mixture $D\in\Delta_{\mathrm{mix}}$ can be represented by a point on the $(K-1)$-dimensional simplex ${\bm{p}}\in\Delta^{K}$. Additionally, as performative response depends on predictor accuracy, the transition operator depends on the classifier $h$ through a $K$-dimensional vector of accuracies $\mathbf{acc}(h)=\left(\mathrm{acc}_{1}(h),\dots,\mathrm{acc}_{K}(h)\right)\in\left[0,1\right]^{K}$. Therefore, transition operators induced by the evolutionary model can be viewed as lower-dimensional mappings ${\bm{p}}^{\prime}=\mathcal{T}({\bm{p}},\mathbf{acc}(h))$ from a $(2K-1)$-dimensional space to a $(K-1)$-dimensional space.
@@
The gaps between the average accuracies of equilibrium points relates to the distinction between performative optimality and performative stability [e.g., 35]. Additionally, convex functions over the simplex (and their sub-gradients) emerge as a key component in our analysis of oracle classifiers (Section 4), and also appears as a key component in the analysis of proper scoring rules and their performative counterparts [36, 6] through the fundamental characterization theorem [15, 44]. From this perspective, interpreting the convex potential of oracle-learner games as a generator of proper scoring rules is an intriguing direction for future work.

### v3

@@ 5 Avenues for Coexistence
Figure 3: Graphical illustration of theoretical coexistence results. Top row presents data distributions, bottom row shows the corresponding prediction games. (Left) Stable and mutualistic coexistence induced by the use of a proxy loss (Theorem˜2). (Center) Stable coexistence induced by interpolation (Section˜D.7). (Right) Mutualistic coexistence induced by a finite training set (Section˜D.8).
@@ 5.1 Learning in Practice
While the goal in learning is to minimize the expected 0-1 loss, in practice learning often resorts to optimizing proxy objectives over sampled data via empirical risk minimization. Here we show that, perhaps surprisingly, basic aspects of this approach—namely the use of surrogate losses, access to finite data, and interpolation methods—can create conditions that facilitate mutualism, in which all groups benefit from coexistence. For each aspect, we demonstrate through a carefully crafted example how it can act as a mechanism that enables stable coexistence. Here we focus our discussion on surrogate loss as a representative construction; for finite data and memorization, we illustrate the underlying distribution and induced game in Figure˜3, and defer the formal analysis to Section˜D.6.
@@
The construction is outlined in Figure˜3 (Left), and proof is provided in Section˜D.6. The proof leverages the bias of hinge loss against minority classes: For each group alone, minimizing the hinge loss biases the learned classifier against the minority class, which is suboptimal. But since each group has a different minority class, these biases cancel out at the mixed equilibrium. The surprising property is that this state is also evolutionarily stable: when one group grows slightly (i.e., $p_{1}=0.5+\epsilon$), it loses more from the fact that the other group shrinks (i.e., $p_{2}=0.5-\epsilon$) than it gains from its own growth. In this sense, each group needs the other to maintain its existence, giving rise to stable and mutualistic symbiosis. Section˜D.6.3 expands on this idea to show that varying the amount of regularization leads to bifurcations, in which the game transitions between 3, 1, and 5 equilibria.
@@
In the example game above, as in those we construct for finite data (Section˜D.8; Figure˜3 (Right)) and interpolation (Section˜D.7; Figure˜3 (Center)), mutualism is enabled by how user self-selection compensates for the algorithm’s imperfections. When the resulting bias is complementary across groups, this can give rise to (implicit) cooperation. One interpretation is that the existence of each group acts as regularization on others: from the perspective of group $k$, the objective can be written as $\operatorname*{arg\,max}_{h}\mathrm{acc}_{k}(h)+\lambda R(h)$ where $R(h)=\sum_{j\neq k}\mathrm{acc}_{k}^{\prime}(h)$ is a data-dependent regularizer with coefficient $\lambda=\frac{1}{p_{k}}\sum_{j\neq k}p_{j}$. Nonetheless, our results should not be taken to imply that proxy losses or finite data are a good means to achieve coexistence; rather, they highlight how mutualism can organically materialize when learning affects different groups differently. We explore this further empirically in Section˜6.2.
@@ 5.2 Stabilizing Coexistence Equilibria
Here we focus on a particular aspect of this general question that arises from our last result in Sec. 4, namely: given a system with a desirable but unstable mixed equilibrium $\bm{p}$$^{*}$, how do we stabilize it? We propose a conceptually simple solution that works by reweighing examples in the standard accuracy objective in a way that accounts for dynamics. The main idea is to invert the natural tendency of dynamics to push away from the unstable equilibrium, and instead pull towards it, by training ‘as if’ the distribution was at a different state $\bm{p}$$^{\prime}$ than its actual current state $\bm{p}$}:
@@
Proof in Section˜D.9. The idea is that $\mathcal{A}^{\prime}$ trains ‘as if’ the state is $2{\bm{p}}^{*}-{\bm{p}}$, which can be achieved by weighing examples from each group $k$ by $w_{k}=\frac{p_{k}}{2p^{*}_{k}-p_{k}}$. This inverts the natural tendency of dynamics to push away from the unstable equilibrium, and instead pull towards it. The statement regards optimal classifiers and requires knowledge of $\bm{p}$$^{*}$, but in practice we can set $w$ according to an estimated $\tilde{{\bm{p}}}$. We further explore this approach in Section˜6.1.
@@ 6 Experiments
To demonstrate how some of the principles we have discussed so far apply in more realistic settings, we now turn to explore evolutionary prediction games empirically using real data and simulated dynamics. We include three experiments. The first experiment complements our results on Sec. 5.2 by demonstrating how stabilization can improve both overall and per-group accuracies via coexistence. The second complements Sec. 5.1, showing how mutualism can arise organically due to group balancing, here from complementarities in label noise. The third experiment explores fairness under population dynamics (Prop. 1) to illustrate the limitations of static fairness measures.
6.1 Mutualistic Coexistence From Data Augmentation
Data augmentation is a method for enriching the training set with class-preserving transformations of inputs (e.g., camera angle, source of lighting) [e.g., 55]. Augmentations are typically considered as artificial constructs, but there are also cases where they emerge naturally (e.g., medical images from different hospitals or imaging devices). Here we show that when different sources correspond to different groups, data pooling can serve to foster beneficial coexistence in which all groups gain from the existence of others. This suggests that ‘natural’ augmentations can be effective also in the long run.
Method.
We use the CIFAR-10 image recognition dataset [36], which consists of 60,000 32x32 color images in 10 classes, with 6,000 images per class. As horizontal image flips are considered class-preserving for this dataset, group $A$ consists of images sampled directly from CIFAR-10, and group $B$ consists of images under horizontal flip. We use a ResNet-9 network for prediction [26], and train it using the ffcv framework with default optimization parameters [38]. For each $\bm{p}$}, we measure the prediction model’s accuracy on the original images from the CIFAR test set (representing $\mathrm{acc}_{A}(h_{\bm{p}})$), and on their flipped counterparts (representing $\mathrm{acc}_{B}(h_{\bm{p}})$). For stabilization, we use the method described in Section˜5.2 with ${\bm{p}}^{*}=(0.5,0.5)$. We simulate the evolutionary process using discrete replicator dynamics, and use linear interpolation to determine intermediate fitness values.
Results.
Figure˜4 (Left) presents the estimated evolutionary prediction game, averaged across 2020{} repetitions. The game has three equilibria: Two stable equilibria with single-group dominance ($92.6\pm 0.1\%{}$ accuracy), and an unstable coexistence equilibrium ($93.5\pm 0.1\%{}$ accuracy). Although fitness is typically higher for the larger group, the mixed equilibrium attains the highest overall population accuracy, suggesting that each groups benefits from the existence of the other. Reaching this point however requires stabilization; Figure˜4 (Center) shows how our algorithm from Section˜5.2 is able to attain stable beneficial coexistence, despite using an empirical proxy objective. Appendix˜F includes additional results on sampling noise, time to convergence, and sensitivity analysis for stabilization.
Figure 4: Empirical evaluation in two-group settings. (Left) Game induced by CIFAR-10 with Resnet-9 and groups representing horizontal flips. The game has unstable mutualism (Section˜6.1). (Center) Simulated stabilization of replicator dynamics on the CIFAR-10 game (Section˜5.2). (Right) Game induced by MNIST with label noise, showing stable mutualism (Section˜6.2).
@@ 6.2 Stable Coexistence From Overparameterization
Modern neural networks are frequently overparameterized, and are known to achieve strong test-time performance despite memorizing the training data [e.g., 64]. Recent research attributes this phenomenon to the implicit biases introduced by the training process [e.g., 60]. Building on these ideas, here we demonstrate that stable coexistence can emerge when an overparameterized neural network is trained on data in which groups have complementing label noise. Our construction considers a multi-class settings where each group includes a majority of noisy examples from some class, and a minority of clean examples from other classes. This aims to captures settings in which user groups (e.g., by country) differ in the marginal distribution $p(y)$ over classes (e.g., spoken language), and where the dominant class is more likely to be misclassified (e.g., due to many accents).
@@
We use the MNIST dataset, set $K=2$, and split the classes unevenly between two groups: Group $A$ is biased towards even digits $\left\{{0,2,4,6,8}\right\}$}, whereas group $B$ is biased towards odd digits, both with a 4:1 imbalance. For each group, we introduce label noise to the majority classes, mapping the true label of each digit $d$ to the next digit with same parity with probability $0.2$ (i.e. for group $A$, label noising stochastically maps $0\mapsto 2$, $2\mapsto 4$, etc.). We train a convolutional neural network for 200 epochs using stochastic gradient decent with momentum. The training and test sets are split independently, and for each $\bm{p}$} we measure the prediction model’s accuracy on both splits of the test sets (representing $\mathrm{acc}_{A}(h_{\bm{p}})$ and $\mathrm{acc}_{B}(h_{\bm{p}})$). Under the given parameters, the label noising process leads to an accuracy upper bound of $84\%$ assuming perfect recognition.
@@ Results.
Figure˜4 (Right) shows the estimated evolutionary prediction game, averaged across 5050{} repetitions. Note how label noise has ’flipped’ the game in that accuracy is higher for the minority group (i.e., $\mathrm{acc}_{B}>\mathrm{acc}_{A}$ when $p_{B}<p_{A}$, and vice versa). This indicates that groups naturally balance each other in this setting. As a result, the game has a stable coexistence equilibrium, with $80.4\pm 0.2\%{}$ test accuracy (vs. the theoretical upper bound of $84\%$ due to label noise). Training accuracy is $98.7\pm 0.1\%{}$, suggesting that interpolation is indeed the likely phenomenon at play. See Section˜D.7 for further discussion on how interpolation facilitates coexistence as an underlying mechanism.
6.3 Population Dynamics and Fairness
Figure 5: Evolutionary dynamics in a three-group setting, induced by a linear SVM trained on the ACSIncome dataset (Section˜6.3). (Left) Replicator dynamics on the three-group simplex. Red line indicates a trajectory starting from the uniform mixture, and background colors indicate basins of attraction. (Center) Population composition over time for the same trajectory. Proportion of NY users becomes negligible at $t\approx 316$. (Right) Fitness over time. Effective group disparity is significant at the outset, but shrinks by an order of magnitude after NY users are driven out.
Lastly, we examine how population dynamics interact with notions of fairness. Our goal here is to show that a system can appear to be fair at a given point in time—but only because some groups have been driven out of the game, and fairness is measured on the remaining observable groups. Our message here is that fairness should consider not only what has happened, but also what could have happened, particularly with respect to groups that perhaps had previously been excluded.
@@ 6.3 Population Dynamics and Fairness
We use a subset of the Folktables ACSIncome dataset [17], a tabular income-level classification task based on US census data. We associate groups with data from three states in the US, with the idea that firms operate in different states but train a model on data aggregated from all sources. Specifically, we set $K=3$, and use 2018 census data from the states of California (195,665 datapoints), New York (103,021), and Texas (135,924). We use linear SVM as the learning algorithm, and randomly hold out 5,000 points from each state for evaluating accuracy. To compute fitness at a given population state $\bm{p}$}, we sample a dataset of size $n=1000$ from the mixture, train a classifier, and measure group accuracies using the held-out sets. Results are averaged over $10$ repetitions to obtain mean fitness values, and we simulate replicator dynamics starting from the uniform population state.
@@ Results.
Figure˜5 shows how evolution transitions between two dynamical phases: an initial phase (up to time step ${\sim}200$) where one group (NY) diminishes while the other two (CA,TX) remain balanced with similar accuracies, and a second phase (from time step ${\sim}300$) in which the two remaining groups compete until one of them dominates. This illustrates a possible shortcoming of static fairness definitions: At the onset, all groups are present, but disparity in accuracy is $\approx 2\%$. As time progresses, the system becomes more "fair" — but only because the first group has been suppressed (i.e., its relative size is $\leq 1\%$). Auditing the prediction algorithm for fairness at this point in time creates an appearance of fairness. This holds only for the two groups that remain (disparity $\approx 0.2\%$), and there is insufficient observed data to help identify the preceding existence of a missing group. Section˜F.4 provides additional support, exploring similar phenomena under different learning algorithms.
@@ Acknowledgements
Appendix A Broader Impact
@@
Scope and limitations.
As the bulk of our results are theoretical in nature, naturally they rely on several assumptions. Some regard the learning process: for example, we consider only simple accuracy-maximizing algorithms that optimize over a fixed-size training set sampled only from the current (and not past) data distribution. Others concern dynamics: for example, we assume no exogenous forces (i.e., which could prevent extinction), intra-group distribution shifts, or inter-group dependencies (other than those indirectly formed through the classifier). Whether our results hold also when these assumptions are relaxed is an important future question.
@@
Diversity vs. homogeneity.
Our results suggests that systems of the type we analyze will tend towards less diverse population compositions. Nonetheless, and as we state throughout, our results should be taken as describing the tendency of population dynamics — when selective pressure is the primary force, when the learning algorithm is optimal retraining, absent any interventions, and at the limiting equilibrium.
Of course, in reality, many systems are diverse. Possible explanations for why a particular system at a given point in time can sustain several groups are:
  • Time: dynamics have not yet reached equilibrium, or converge very slowly.
  • Learning: the algorithm is not optimal (e.g., uses a proxy or finite data), or is designed to balance groups (e.g., via steering, regularization, or robust training).
  • Intervention: system administrators may intentionally intervene to facilitate diversity; for example, consider how firms invest significant resources in marketing (e.g., Meta reports investing about \$2B per year [^2] Meta Platforms, Form 10-K - Annual report, 2024, SEC Accession No. 0001326801-25-000017. “Advertising Expense”.) to ensure a constant influx of users from diverse social groups.
  • Dynamics: Other forces are at play, such as lack of alternatives, cross-group influence, rigidity of demand, non-linear heterogeneous utility, etc.
Nonetheless, an important but subtle point is that when we observe a system consisting of several groups, we do not observe other groups that could have been supported, as they may have already been driven out. Thus, what we perceive as a mixed state may actually be a snapshot of a process in which “weaker” groups have already vanished, with others possibly to follow.
In biology, the question of coexistence is prevalent, despite consensus that natural selection is the primary force at play. The study of coexistence therefore focuses on how it is sustained despite natural selection. This has revealed many biological and ecological mechanisms that can facilitate coexistence, such as resource partitioning, environment variation, and stabilization via intermediates. In this work we focus on learning itself as a possible stabilizing mechanism (Sections 5,6), which we view as a first step towards the study of coexistence in this context (as also noted by Reviewer oLKx). We also believe that prediction games in particular, and evolutionary game theory more generally, can aid future research in exposing additional mechanisms that can facilitate diversity in environments with learning.
Practical implications.
Evolutionary prediction games can help designers and operators of ML systems anticipate performative population shifts, and integrate this foresight into their decision-making. One key message is that applying conventional retraining risks favoring some groups over others in the long term, even if unintentionally. Evolutionary prediction games can help understand which groups are at risk, why they are at risk, and highlight possible avenues for interventions, if such are needed. One example is to decide on using a stabilizing learning algorithm (such as the one we propose in Prop. 2) instead of retraining to help steer towards desired states. Another is to target the retention of at-risk groups, either by providing benefits or reducing costs (e.g., via subsidies), in order to improve their fitness. A third is to understand if and when to increase the influx of certain groups, e.g. via marketing efforts, to balance the natural tendencies of the system. We will extend the discussion to emphasize these insights.
@@ B.1 Performative Prediction
Performative prediction [48] is a framework for modeling the effects of predictive models on the data distributions they are tested and trained on (see [24] for a recent survey). Dynamics induced by evolutionary prediction games are a special case of the stateful performative prediction framework [5, 50], which models stateful feedback loops. The performative effects of a predictor $h\in\mathcal{H}$ are formally modeled using a transition map $D_{t+1}=\mathcal{T}(D_{t},h)$, where $D_{t},D_{t+1}\in{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$ are the data distributions at times $t,t+1$, and $h\in\mathcal{H}$ is the predictor being deployed. The fixed points of the dynamics induced by evolutionary prediction games (i.e. equilibrium states $\bm{p}$$^{*}$), correspond to fixed point distributions as defined in [5, Definition 4], as they satisfy $D_{{\bm{p}}^{*}}=\mathcal{T}(D_{{\bm{p}}^{*}},h_{{\bm{p}}^{*}})$, where $h_{{\bm{p}}^{*}}$ is the classifier trained on $D_{{\bm{p}}^{*}}$. Moreover, we note that the notion of evolutionary stability is stronger than the notion of fixed point distributions, as evolutionarily stable points also remain stable under small perturbations around $\bm{p}$$^{*}$.
@@
Formally, given a collection of group distributions $D_{1},\dots,D_{K}\in{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$, denote the space of mixtures by $\Delta_{\mathrm{mix}}={\left\{{\sum_{k}p_{k}D_{k}\mid{\bm{p}}\in\Delta^{K}}\right\}}$, and note that $\Delta_{\mathrm{mix}}\subseteq{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$. Any mixture $D\in\Delta_{\mathrm{mix}}$ can be represented by a point on the (K1)(K-1)-dimensional simplex ${\bm{p}}\in\Delta^{K}$. Additionally, as performative response depends on predictor accuracy, the transition operator depends on the classifier $h$ through a $K$-dimensional vector of accuracies $\mathbf{acc}(h)=\left(\mathrm{acc}_{1}(h),\dots,\mathrm{acc}_{K}(h)\right)\in\left[0,1\right]^{K}$. Therefore, transition operators induced by the evolutionary model can be viewed as lower-dimensional mappings ${\bm{p}}^{\prime}=\mathcal{T}({\bm{p}},\mathbf{acc}(h))$ from a (2K1)(2K-1)-dimensional space to a (K1)(K-1)-dimensional space.
@@
The gaps between the average accuracies of equilibrium points relates to the distinction between performative optimality and performative stability [e.g., 43]. Additionally, convex functions over the simplex (and their sub-gradients) emerge as a key component in our analysis of oracle classifiers (Section˜4), and also appears as a key component in the analysis of proper scoring rules and their performative counterparts [45, 8] through the fundamental characterization theorem [21, 53]. From this perspective, interpreting the convex potential of oracle-learner games as a generator of proper scoring rules is an intriguing direction for future work.